# Trapped-ion quantum simulation of the Fermi-Hubbard model as a lattice gauge theory using hardware-aware native gates

### v1

@@ Abstract
The Fermi-Hubbard model (FHM) is a simple yet rich model of strongly interacting electrons with complex dynamics and a variety of emerging quantum phases. These properties make it a compelling target for digital quantum simulation. Trotterization-based quantum simulations have shown promise, but implementations on current hardware are limited by noise, necessitating error mitigation techniques like circuit optimization and post-selection. A mapping of the FHM to a $\mathbb{Z}_{2}$ was recently proposed that restricts the dynamics to a subspace protected by additional symmetries, and its ability for post-selection error mitigation was verified through noisy classical simulations. In this work, we propose and demonstrate a suite of algorithm-hardware co-design strategies on a trapped-ion quantum computer, targeting two key aspects of NISQ-era quantum simulation: circuit compilation and error mitigation. In particular, a novel combination of iteratively preconditioned gradient-descent (IPG) and subsystem von Neumann Entropy compression reduces the 2-qubit gate count of FHM quantum simulation by 35%, consequently doubling the number of simulatable Trotter steps when used in tandem with error mitigation based on conserved symmetries, debiasing and sharpening techniques. Our work demonstrates the value of algorithm-hardware co-design to operate digital quantum simulators at the threshold of maximum circuit depths allowed by current hardware, and is broadly generalizable to strongly correlated systems in quantum chemistry and materials science.
@@ I Introduction
The Fermi-Hubbard Model (FHM) is a simple yet powerful model that brings insight into materials’ electronic and magnetic properties [1, 2]. It is promising in studies of quantum magnetism, antiferromagnetism, high-temperature superfluidity, d-wave pairing and superconductivity, and transitions between these different regimes of the rich phase diagram [3, 4, 5, 6]. Ultracold atoms, trapped ions and semiconductor quantum dots have shown significant experimental progress in analog quantum simulation of the FHM [7, 8, 9, 10, 11, 12]; however, digital quantum simulation on quantum computers remains challenging due to the infidelities in qubits and gate operations on today’s noisy intermediate-scale quantum (NISQ) devices. A lattice gauge theory (LGT) approach for digital quantum simulation of the one-dimensional FHM has been recently proposed and analyzed [13]. Although requiring more qubits and deeper circuits than a direct encoding, the LGT approach allows for a set of local operators on the links between the lattice sites to be fixed to eigenvalues of ±$\pm 1$ depending on the sign of the interaction term in the Hubbard Model. It is typical for LGTs to introduce additional conservation laws that restrict time evolution to a small sector of the entire Hilbert space. For example, in quantum link models for lattice quantum electrodynamics [14, 15, 16, 17, 18, 19, 20, 21], time evolution is naturally or artificially restricted to the sector that respects Gauss’s law, and this has been used for post-selection error mitigation [22, 23, 24, 25]. In a similar spirit, the $\mathbb{Z}_{2}$ encoding of the FHM, in addition to global spin conservation, introduces a large number of local conserved charges [13], which can enable efficient post-selection error mitigation.
@@
More generally, today’s NISQ era digital quantum computers, although large enough to conduct nontrivial gate operations, are limited by qubit coherence time, larger-than-ideal gate execution time, noise in operations, and differing qubit basis sets dependent on the qubit platform. Therefore, minimizing the depth of circuits – often achieved through transpilation, is critical in modeling challenging systems such as the Fermi-Hubbard model.
@@
Despite the benefit of an additional post-selection error mitigation technique introduced by the aforementioned LGT encoding, three significant challenges must be surmounted for Trotterized simulation of the FHM encoded in this fashion on a currently available quantum processing unit (QPU). First, transpiling Trotter steps made up of gates non-native to a QPU can significantly increase overall circuit depth, possibly rendering a circuit unfeasible to execute. This challenge is combated with gradient-based circuit optimization. However, this leads to our second challenge: Classically parameterizing and optimizing a Trotter step is infeasible beyond a small number of qubits, due to the Hilbert space of a circuit increasing exponentially with the number of qubits. Such a problem is ubiquitous in variational algorithms [26, 27, 28, 29], including calculating ground state energies on a variational quantum eigensolver [30, 31, 32, 33]. Finally, ensuring that an ansatz possesses the expressibility needed to represent a desired unitary requires optimizing over many parameters, which is difficult for commonly used gradient-based optimizers due to the presence of local minima, sensitivity to noise, and sensitivity to initial starting conditions.
@@
Here, we surmount these challenges by exploiting the symmetry of an objective unitary to allow for classically feasible circuit optimization. Specifically, we use the Trotter step of a Fermi-Hubbard Model via $\mathbb{Z}_{2}$ transformation as our objective unitary and decompose this unitary as a combination of repeating, static three-qubit circuits. We then create a hardware-aware, parameterizable, three-qubit ansatz using the native gateset of the IonQ Aria QPU, and optimize this hardware-aware ansatz using a newly developed optimization technique –- that of iteratively pre-conditioned gradient descent (IPG) [34] –- and compare the effectiveness of IPG to that of commonly leveraged optimizers for gradient-based circuit optimization. Finally, we run the circuits on the IonQ Aria QPU, leveraging debiasing and sharpening as error mitigation techniques combined with the conserved-symmetry-based error mitigation approach inherent to lattice gauge theories. This novel, systematic method of approaching circuit optimization on a Trotterized system results a 14% reduction in the number of native two-qubit Mølmer-Sørenson (MS) gates per Trotter step, combined with an even larger reduction in the number of single qubit gates. A further reduction of 25% is effected by using von Neumann entropy based arguments for reducing circuit depth while maintaining sufficient expressibility, resulting in a total 37% reduction in 2-qubit gate count. This, combined with debiasing in tandem with symmetry-based error mitigation, demonstrates a strategy to reliably run a large depth circuit of a fermionic system on present day QPU hardware.
@@
The structure of the paper is the following. In Sec. II we introduce and outline the quantum circuit of the Trotterized $\mathbb{Z}_{2}$ Fermi Hubbard Model. In Sec. III we decompose the circuit into three-qubit sub-circuits and create a thirty-parameter, hardware-aware ansatz with adequate expressibility. In Sec. IV we introduce iteratively pre-conditioned gradient descent (IPG) in the context of our specific quantum circuit optimization. In Sec. V we optimize the sub-circuit with IPG, and combine sub-circuits to create an overall optimized Trotter step. In sec. VI we introduce the several error mitigation techniques we employed. In Sec. VII we run multiple trotter steps on the IonQ Aria QPU with Debiasing and Sharpening as active error mitigation strategies. Sec. VIII reports results on further reducing two-qubit gate counts using von Neumann entropy arguments while maintaining sufficient expressibility to represent the desired quantum circuit. We conclude in Sec. IX.
@@ II.1 Fermi Hubbard Model
The standard way to map the FHM on a universal gate-based quantum computer involves the choice of a specific spin representation combined with a Jordan-Wigner transformation to convert the fermion operators to qubits [37, 13, 38]. Here we choose the so-called slave-spin representation and a 1D lattice for the subsequent Jordan-Wigner transformation since that preserves the nearest-neighbor (local) nature of the couplings in the Hamiltonian of Eq. (1) [39, 40]. Note that higher-dimensional FHM lattices require long-range nonlocal couplings, which can be implemented in the trapped-ion QC architecture we discuss in Secs. III.2 and VII since this QC hardware platform possesses all-to-all qubit connectivity without requiring SWAP gates. Additionally, this means that the geometry of a lattice does not need to be mapped to a physical qubit layout, as is required in QC architectures possessing solely next-neighbor connectivity [41, 42, 43].
@@
with the pseudospins $I_{i}$ satisfying the standard SU(2) Lie algebra $[I_{i},I_{j}]=i\epsilon_{ijk}I_{k}$, with $\epsilon_{ijk}$ being the fully antisymmetric Levi-Civita symbol. One route to implementing this on a quantum computer is to apply the Jordan-Wigner transformation directly to this encoding, thus requiring $2N$ qubits to represent $N$ sites [39]. Another route is to transform the FHM Hamiltonian into the form of a lattice gauge theory, which we discuss next. This route is incentivized by using conserved additional local and global symmetries for error mitigation and post selection, described in the ensuing Sections III.1, VI.2. Note that the introduction of auxiliary pseudospins enlarges the Hilbert space beyond that of the original FHM [13], and hence accompanying constraints must be applied to restrict the simulation to the physically relevant subspace of the enlarged Hilbert space.
II.2 $\mathbb{Z}_{2}$ Transformation
@@
Using such a transformation, and assuming translational symmetry $J_{ij,\sigma}=J$, the $\mathbb{Z}_{2}$ Hamiltonian reads,
@@
The major change that engenders this to be a $\mathbb{Z}_{2}$ is the mediation of the hopping between neighboring sites by a dynamical spin operator $\hat{\tau}$ with eigenvalues ±$\pm 1/2$, instead of a number $J_{ij}$.
@@ II.3 Mapping fermions to spins: Jordan-Wigner Transformation
The number operators on each site $\hat{n}_{i\sigma}=\hat{c}_{i\sigma}^{\dagger}\hat{c}_{i\sigma}$ have eigenvalues $\{0,1\}$ whereas the spin operators have eigenvalues ±$\pm 1/2$. They are related by $\hat{S}_{i\sigma}^{z}=\hat{c}_{i\sigma}^{\dagger}\hat{c}_{i\sigma}-1/2$. Note that in our encoding, the occupied fermionic state is represented by $|0\rangle$ and the unoccupied state is represented by $|1\rangle$ in the computational basis of the qubits, and these have eigenvalues $+1$ and $-1$ respectively when measured in the $Z$ basis. Thus, the explicit relationship between the quantum computer expectation values reads,
@@ II.4 Trotterization of LGT Hamiltonian
$H_{\rm LGT}=$displaystyle-- $4J\sum_{i,\sigma}\left(\hat{\tau}^{z}_{i,i+1}\hat{S}^{+}_{i\sigma}\hat{S}^{-}_{i+1,\sigma}+{\rm H.c.}\right)$ (7)
@@
II.5 Direct quantum circuit of a single Trotter step
@@ III Parameterized single trotter step ansatz
Figure 1: Explicit circuit decomposition of $\hat{C}$ and $\hat{B}$ for a single Trotter step of $H_{LGT}$ from [13]. $\theta_{1}=2J\Delta t$, $\theta_{2}=-2J\Delta t$, and $\theta_{3}=U\Delta t$
@@ III.1 Mapping FHM Hamiltonian’s symmetries to a modular circuit ansatz
To construct a Trotter step, we used the rules of fermionic encoding via Jordan Wigner and the properties of a $\mathbb{Z}_{2}$ to build a time evolution of $H_{LGT}$ in terms of permutations of $\hat{C}$ and $\hat{B}$.
A subroutine that is easily simplifiable contains as little number of qubits and gates as needed, while appearing enough times in a circuit to justify optimization. Since $\Delta t$ remains constant during $H_{LGT}$ time evolution, $\hat{C}$ and $\hat{B}$ are identical regardless of the number of trotter steps. Therefore, a separate combination of gates equivalent to $\hat{C}$ can be substituted in place of its original gate decomposition to reconstruct the overall Trotter step of $H_{LGT}$.
@@
Each permutation of $\hat{C}$ only acts on one three-qubit subspace at a time, making it easy to classically simulate and optimize over. Furthermore, $\hat{C}$ occurs $2N$ times per Trotter step, where $N$ is the number of sites, compunding any improvement to $\hat{C}$ across the entire circuit. This is especially true in the context of single and two-qubit gates: Two-qubit gate errors can exhibit one to three orders of magnitude higher infidelity than single qubit gates, even amongst recent state-of-the-art reports [47] [48], resulting in two-qubit gate errors dominating single-qubit errors at high gate counts. Therefore, any two-qubit gate reduction in $\hat{C}$ can increase the number of Trotter steps that can be reliably executed on a NISQ device. Section VIII introduces a more systematic way to more aggressively decrease the two-qubit gatecount per Trotter step.
@@
By imposing cyclic boundary conditions on the FHM, $\hat{C}$ will operate on qubits that are physically distant on a chip, which is compatible with IonQ’s trapped-ion systems without extra SWAP gates due to all-to-all qubit connectivity.
@@
Finally, quantum circuits submitted to cloud accessible QPUs undergo transpilation. Therefore, it is beneficial to simplify a circuit into a gateset that is easily mappable to the QPU in question. Here the concept of a native gateset is important, which we discuss next.
III.2 Hardware constraints of a trapped-ion quantum processor
III.2.1 Trapped-Ion Native Gateset
When a circuit representation of a quantum computation is submitted to an IonQ QPU via the default API, it goes through a compilation pipeline. First, the circuit will be transpiled into a standard universal gate set. Then it is optimized using a quantum circuit optimizer (QCO) specifically designed for such a gate set. After the optimization, the circuit is converted to native gates and executed. In this sense, native gates are the unitary transformations implemented directly as tailored laser pulses acting on the ions, without further gate-based decomposition.
IonQ’s native gates API allows direct access to two single-qubit gates and various two-qubit entangling gates depending on the specific system. The two single-qubit gates are referred to as $GPi(\phi)$ and $GPi2(\phi)$ :
$GPI(\phi)=\begin{pmatrix}0&e^{-i\phi}\\ e^{i\phi}&0\end{pmatrix}$ (11)
$GPI2(\phi)=\frac{1}{\sqrt{2}}\begin{pmatrix}1&-ie^{-i\phi}\\ -ie^{i\phi}&1\end{pmatrix}$ (12)
These are $\pi$ and $\pi/2$ rotations of the qubit along an axis that is specified by $\phi$ on the equatorial plane of the Bloch sphere [49].
On the 32-qubit Aria systems, the supported native two-qubit gate is referred to as the arbitrary angle Molmer-Sorensen (MS) Gate [49]. This is a simultaneous rotation of the target qubit 0 and target qubit 1 along the axis specified by $\phi_{0}$ and $\phi_{1}$ on the equatorial plane of the Bloch sphere, respectively. With full control over the provided degrees of freedom, these three gates make a universal gate set.
Native gates are popular for a few reasons. In quantum characterization, verification, and validation (QCVV) studies, the native gates API allows bypassing optimizations. This is crucial, for example, to echo-type experiments that would otherwise be interrupted by optimization.
The native gates also provide a pathway to improve the efficacy of circuit executions. For example, a $ZZ(\theta)$ type gate, which is by default implemented as two maximally entangling $MS(\pi/2)$ gates and one $R_{z}(\theta)$ rotation, can be implemented using just a single $MS(\theta)$ gate. This leads to a non-trivial improvement since entangling gates are an important source of noise. We note that for typical implementations of MS gates, the magnitude of noise is positively correlated with the gate angle. For example, a study using one of the now-decommissioned system [50] reported improvement from  97.5% to  99.6% two-qubit fidelity when comparing gate angles $\theta$’s of $\pi/2$ and $\pi/100$, respectively. This makes native gates particularly preferable when a non-negligible small-angle entangling gate is required for the computation.
III.3 Determining simplifiable subroutines

### v2

@@ Abstract
The Fermi-Hubbard model (FHM) is a simple yet rich model of strongly interacting electrons with complex dynamics and a variety of emerging quantum phases. These properties make it a compelling target for digital quantum simulation. While Trotterization-based quantum simulations have shown promise, implementations on current hardware are limited by noise, necessitating error mitigation techniques like circuit optimization and post-selection. A mapping of the FHM to a $\mathbb{Z}_{2}\$LGT was recently proposed that restricts the dynamics to a subspace protected by additional symmetries, but a faithful representation of its dynamics on quantum hardware remains a challenge. In this work, we propose and demonstrate a suite of algorithm-hardware co-design strategies on a trapped-ion quantum computer, targeting two key aspects of NISQ-era quantum simulation: circuit compilation and error mitigation. To this end, we apply iteratively preconditioned gradient descent (IPG), which reduces the 2-qubit gate count of the FHM simulation by 14%. When combined with von Neumann entropy (VNE) compression, this results in a total reduction of 36%. We observe in QPU experiments that IPG alone, with error mitigation based on conserved symmetries (debiasing and sharpening techniques), enables a faithful observation of domain wall dynamics —even in the presence of noise that attenuates the magnitude of these overall dynamics. Using IPG in tandem with VNE compression in a noisy classical statevector simulator, we faithfully double the number of simulatable Trotter steps. Our work demonstrates the value of algorithm-hardware co-design to operate digital quantum simulators at the threshold of maximum circuit depths allowed by current hardware, and is broadly generalizable to strongly correlated systems in quantum chemistry and materials science.
@@ I Introduction
The Fermi-Hubbard Model (FHM) is a simple yet powerful model that brings insight into materials’ electronic and magnetic properties [1, 2]. It is promising in studies of quantum magnetism, antiferromagnetism, high-temperature superfluidity, d-wave pairing and superconductivity, and transitions between these different regimes of the rich phase diagram [3, 4, 5, 6]. Although analog quantum simulation based on ultracold-atom, trapped-ion and semiconductor-quantum-dot experiments have shown significant progress in quantum simulation of the FHM [7, 8, 9, 10, 11, 12], such explorations are typically limited to regions defined by the experiment setup. In contrast, digital quantum simulation is universal in nature, but in practice challenging with today’s noisy intermediate-scale quantum (NISQ) devices. There are many paths to address these challenges.
@@
It is possible to encode the simulation into a symmetry protected subspace. A lattice gauge theory (LGT) approach for digital quantum simulation of the one-dimensional FHM has been recently proposed and analyzed [13]. Although requiring more qubits and deeper circuits than a direct encoding, the LGT approach allows for a set of local operators on the links between the lattice sites to be fixed to eigenvalues of $\pm 1$ depending on the sign of the interaction term in the Hubbard Model. It is typical for LGTs to introduce additional conservation laws that restrict time evolution to a small sector of the entire Hilbert space [14]. For example, in quantum link models for lattice quantum electrodynamics [15, 16, 17, 18, 19, 20, 21, 22], time evolution is naturally or artificially restricted to the sector that respects Gauss’s law, and this has been used for post-selection error mitigation [23, 24, 25, 26]. In a similar spirit, the $\mathbb{Z}_{2}\$LGT encoding of the FHM, in addition to global spin conservation, introduces a large number of local conserved charges [13], which can enable efficient post-selection error mitigation.
@@
Circuit compilation is another effective technique. More generally, today’s NISQ era digital quantum computers, although large enough to conduct nontrivial gate operations, are limited by qubit coherence time, larger-than-ideal gate execution time, noise in operations, and differing qubit basis sets dependent on the qubit platform. Therefore, minimizing the depth of transpiled circuits is critical in modeling challenging systems such as the Fermi-Hubbard model.
@@
Directly executing Trotterized FHM simulations encoded as a $\mathbb{Z}_{2}\$LGT on current quantum processing units (QPUs) remains infeasible without first optimizing each Trotter step. A parametrized circuit template, known as an ansatz, can be used to optimize each Trotter step. However, classically parameterizing and optimizing a Trotter step ansatz is infeasible beyond a small number of qubits, due to the Hilbert space of a circuit increasing exponentially with the number of qubits. Such a problem is ubiquitous in variational algorithms [27, 28, 29, 30], including calculating ground state energies on a variational quantum eigensolver [31, 32, 33, 34]. Even if a Trotter step can be classically optimized, ensuring that an ansatz possesses the expressibility needed to represent it requires attempting different numbers of single- and 2-qubit gates and their arrangements, and optimizing each gate angle. With a large number of parameters and no guarantee that a specific gate arrangement is sufficiently expressible, commonly used gradient-based optimizers struggle to converge to a solution (this is elaborated on in Sec. IV.3). Additionally, transpiling Trotter steps made up of gates non-native to a QPU can significantly increase overall gate count, further rendering a circuit infeasible to execute.
@@
Here, we address these challenges by exploiting the symmetry of an objective unitary to allow for classically feasible circuit optimization. Specifically, we use the Trotter step of a Fermi-Hubbard Model via $\mathbb{Z}_{2}\$LGT transformation as our target unitary and decompose this unitary as a combination of repeating, static three-qubit circuits. We then create a hardware-aware, parameterizable, three-qubit ansatz using the native gateset of the IonQ Aria QPU, and optimize this ansatz using a newly developed optimization technique –- that of iteratively pre-conditioned gradient descent (IPG) [35] –- and compare the effectiveness of IPG to that of commonly used gradient-based optimizers. Finally, we run the circuits on the IonQ Aria QPU, leveraging debiasing and sharpening as error mitigation techniques combined with the conserved-symmetry-based error mitigation approach inherent to lattice gauge theories.
This novel, systematic method of approaching circuit optimization on a Trotterized system results a 14% reduction in the number of native two-qubit gates (Appendix, Sec. VIII.1) per Trotter step, combined with an even larger reduction in the number of single qubit gates. A further reduction of 22% is effected by using von Neumann entropy based arguments for reducing circuit depth while maintaining sufficient expressibility, resulting in a total 36% reduction in 2-qubit gate count. This, combined with debiasing and symmetry-based error mitigations, demonstrates a strategy to reliably run a large depth circuit of a fermionic system on present day QPU hardware.
@@
The structure of the paper is the following. In Sec. II we introduce and outline the quantum circuit of the Trotterized $\mathbb{Z}_{2}\$LGT Fermi Hubbard Model. In Sec. III we decompose the circuit into three-qubit sub-circuits and create a thirty-parameter, hardware-aware ansatz with adequate expressibility. In Sec. VIII.2 we introduce iteratively pre-conditioned gradient descent (IPG) in the context of our specific quantum circuit optimization. In Sec. IV.3 we optimize the sub-circuit with IPG, and combine sub-circuits to create an overall optimized Trotter step. In Sec. IV.4 we create a further optimized circuit via von Neumann entropy compression. In sec. V we introduce the several error mitigation techniques we employed. In Sec. VI.2 we run multiple trotter steps on the IonQ Aria QPU with Debiasing and Sharpening as active error mitigation strategies, but without VNE compression. In Sec. VI.3 we run multiple trotter steps optimized with VNE compression on a depolarizing noise statevector simulator. We conclude in Sec. VII.
@@ II.1 Fermi Hubbard Model
The standard way to map the FHM on a universal gate-based quantum computer involves the choice of a specific spin representation combined with a Jordan-Wigner transformation to convert the fermion operators to qubits [38, 13, 39]. Here we choose the so-called slave-spin representation and a 1D lattice for the subsequent Jordan-Wigner transformation since that preserves the nearest-neighbor (local) nature of the couplings in the Hamiltonian of Eq. (1) [40, 41]. Note that higher-dimensional FHM lattices require long-range nonlocal couplings, which can be implemented in the trapped-ion QC architecture we discuss in Secs. VIII.1 since this QC hardware platform possesses all-to-all qubit connectivity without requiring SWAP gates. Additionally, this means that the geometry of a lattice does not need to be mapped to a physical qubit layout, as is required in QC architectures possessing solely next-neighbor connectivity [42, 43, 44].
@@
with the pseudospins $I_{i}$ satisfying the standard SU(2) Lie algebra $I_{i},I_{j}]=i\epsilon_{ijk}I_{k}$, with $\epsilon_{ijk}$ being the fully antisymmetric Levi-Civita symbol. One route to implementing this on a quantum computer is to apply the Jordan-Wigner transformation directly to this encoding, thus requiring $2N$ qubits to represent $N$ sites [40]. Another route is to transform the FHM Hamiltonian into the form of a lattice gauge theory, which we discuss next. This route is incentivized by using conserved additional local and global symmetries for error mitigation and post selection, described in the ensuing Sections III.1, V.2. Note that the introduction of auxiliary pseudospins enlarges the Hilbert space beyond that of the original FHM [13], and hence accompanying constraints must be applied to restrict the simulation to the physically relevant subspace of the enlarged Hilbert space.
II.2 $\mathbb{Z}_{2}\$LGT Transformation
@@
Using such a transformation, and assuming translational symmetry $J_{ij,\sigma}=J$, the $\mathbb{Z}_{2}\$LGT Hamiltonian reads,
@@
The major change that engenders this to be a $\mathbb{Z}_{2}\$LGT is the mediation of the hopping between neighboring sites by a dynamical spin operator $\hat{\tau}$ with eigenvalues $\pm 1/2$, instead of a number $J_{ij}$.
@@ II.3 Mapping fermions to spins: Jordan-Wigner Transformation
The number operators on each site $\hat{n}_{i\sigma}=\hat{c}_{i\sigma}^{\dagger}\hat{c}_{i\sigma}$ have eigenvalues $\{0,1\}$ whereas the spin operators have eigenvalues $\pm 1/2$. They are related by $\hat{S}_{i\sigma}^{z}=\hat{c}_{i\sigma}^{\dagger}\hat{c}_{i\sigma}-1/2$. Note that in our encoding, the occupied fermionic state is represented by $|0\rangle$ and the unoccupied state is represented by $|1\rangle$ in the computational basis of the qubits, and these have eigenvalues +1+1 and 1-1 respectively when measured in the $Z$ basis. Thus, the explicit relationship between the quantum computer expectation values reads,
@@ II.4 Trotterization of LGT Hamiltonian
$H_{\rm LGT}=$ - $4J\sum_{i,\sigma}\left(\hat{\tau}^{z}_{i,i+1}\hat{S}^{+}_{i\sigma}\hat{S}^{-}_{i+1,\sigma}+{\rm H.c.}\right)$ (7)
@@
III Parameterized single trotter step ansatz
@@
Figure 1: Direct circuit decomposition of $\hat{C}$ and $\hat{B}$ for a single Trotter step of $H_{LGT}$ from [13]. $\theta_{1}=2J\Delta t$, $\theta_{2}=-2J\Delta t$, and $\theta_{3}=U\Delta t$
@@ III.1 Mapping FHM Hamiltonian’s symmetries to a modular circuit ansatz
Keeping the timestep $\Delta t$ during trotterized time-evolution of a system is recommended on a digital quantum computer. This means that each permutation of $\hat{C}$ only acts on one three-qubit subspace at a time, allowing for straightforward classical simulation of $\hat{C}$ and efficient reconstruction of each overall trotter step. Additionally, $\hat{C}$ occurs $2N$ times per Trotter step (where $N$ is the number of sites), compounding any improvement to $\hat{C}$ across an entire trotter step. This is especially true in the context of single and two-qubit gates: Two-qubit gate errors can exhibit one to three orders of magnitude higher infidelity than single qubit gates, even amongst recent state-of-the-art reports [48] [49], resulting in two-qubit gate errors dominating single-qubit errors at high gate counts. Therefore, any two-qubit gate reduction in $\hat{C}$ can increase the number of Trotter steps that can be reliably executed on a NISQ device. Section IV.4 introduces a more systematic way to more aggressively decrease the two-qubit gatecount per Trotter step.
@@
Imposing cyclic boundary conditions on the Fermi-Hubbard Model requires $\hat{C}$ to act on physically distant qubits on a chip. This is achievable on IonQ’s trapped-ion systems without extra SWAP gates due to all-to-all qubit connectivity.
@@
Finally, quantum circuits submitted to cloud accessible QPUs undergo transpilation. Therefore, it is beneficial to simplify a circuit into a gateset that is easily mappable to the QPU in question. Here the concept of a native gateset is important, which we discuss in the Appendix, Sec. VIII.1.
IV Algorithm-Hardware Codesign of the Circuit
In this section, we detail our approach that reduces the two-qubit gate count of the Trotterized $\mathbb{Z}_{2}\$LGT Fermi–Hubbard model by up to 36%.
IV.1 Determining simplifiable subroutines